OpenStudy (okdutchman7):

Please help. WILL MEDAL! @Mehek14 @triciaal @sleepyjess @hartnn @satellite73 @jim_thompson5910

OpenStudy (okdutchman7):

\[\sin \theta = -\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (erak):

Do you have a restriction on theta?

OpenStudy (okdutchman7):

\[0 \le \theta < 360\]

OpenStudy (okdutchman7):

I think 300 is a solution, but I have to find them all. There is supposedly 2.

OpenStudy (anthonyym):



OpenStudy (erak):

Were you taught the CAST rule and something called the 30-60-90 triangle?

OpenStudy (okdutchman7):

Not the CAST rule but yes I am aware that this forms a 30-60-90 triangle.

jimthompson5910 (jim_thompson5910):

One method is to use the unit circle https://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/2000px-Unit_circle_angles_color.svg.png Look for points (on the unit circle) that have y coordinates equal to \(\LARGE -\frac{\sqrt{3}}{2}\)

OpenStudy (okdutchman7):

So the second value would be 240.

OpenStudy (okdutchman7):

Thanks for that resource @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anthonyym):

 ASTC All: sin, cos, tan are all positive S: only sin positive T: only tan positive C: only cos positive Therefore sine is negative in quadrants III and IV, which matches -sqrt3 / 2, which is negative.